# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@ 4 Applications
In this section, we apply the main results to multifractal analysis and mean Li-Yorke chaos. The main results are Theorems 4.1 and 4.3.
@@ 4.1 Variational principle of level sets
In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ in the sense of an abstract framework.
Let $(X,f)$ be a dynamical system, and let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function. Let
$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right],\$
For $\alpha\in L$, let $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$, and let
$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\}$
denote the level set of $\alpha$ w.r.t. $\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$. Then the continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
Condition A: Given a continuous function $\varphi$ on $M_{f}(X)$, for every $\alpha\in L_{\varphi}$ we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
Some examples of $\varphi$ satisfying Condition A are given as follows:
@@
    Let $\psi$, $\phi\in C(X)$ such that $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ holds for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$). We define
    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu}.$
@@
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i.e. for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
@@
    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for any $\Phi$ and $\alpha\in L_{\Phi}$.
    Specially, when the sequence $\{\phi_{n}\}$ is an additive function $\{S_{n}\psi\}$ with $\psi\in C(X)$, in this case we have
@@
Given a nonempty open subset $U\subset X$, we put
@@
Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset. Suppose that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A. Then for any $\alpha\in L$ and $S\in\{B,P\}$, we have
@@
Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$, which is not dependent on $\mu$, such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$ (4.1)
holds for all $\mu\in M_{f}(X)$ and any small enough $\epsilon>0$. Together with the fact $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
@@
for all sufficiently small $\epsilon>0$, which implies that
@@
Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, then $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
@@
A direct consequence of Theorem 4.1 is the variational principles for Bowen metric mean dimension of level sets obtained by Backes and Rodrigues [BR23, Theorem A], but we can strength the result which states that the variational principle is still valid for packing metric mean dimension of level sets.
@@
Let $(X,d,f)$ be a dynamical system with specification property and $\psi$ be a continuous function on $X$. Let
@@
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
@@
Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We get the desired result by Theorem 4.1. ∎
Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure. Then, by Theorem 4.1, one can get analogous variational principles for the metric mean dimensions of level sets of some special functions, which are determined by the dynamical behaviors of the system.
@@ 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy. We give a quantitative characterization for the chaotic phenomenon of infinite entropy systems. Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
Let $(X,d,f)$ be a dynamical system. This naturally gives a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is compact endowed with the product metric $d\times d$ defined by $d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\}$, and the map
$f\times f:X\times X\rightarrow X\times X$
given by $(f\times f)(x,y):=(fx,fy)$ is continuous.
It is well-known that if two systems admit the specification property, then the product system also has the specification property(cf. [DGS76, Proposition 21.4]). A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
The set of mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$. A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set.
@@
Let $(X,d,f)$ be a dynamical system with specification property. Then
$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d),$
$\overline{\rm mdim}_{M}^{UC}(f\times f,{\rm MLY}(f),d\times d)=\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d).$
@@
Claim: there exists $\mu_{0}\in M_{f\times f}(X\times X)$ with $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
Proof of the claim.
Specification property implies that the system has positive entropy. Then by the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$. We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then any an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim. ∎
Now take $\nu\in M_{f}(X)$ and put $\mu_{1}=\phi_{\ast}\nu\in M_{f\times f}(X\times X)$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Assume that $\overline{\rm mdim}_{M}(f,X,d)$ is finite. For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$ (4.4)
Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then for every $k$,
$\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)=\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta).$ (4.5)
Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ so that ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
$H_{\rho_{k}^{\ast}}(\eta^{n})$ $=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=-\sum_{C\in\xi^{n}}\rho_{k}^{\ast}(C\times C)\log\rho_{k}^{\ast}(C\times C)$
$=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=H_{\rho_{k}}(\xi^{n}),$
which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\leq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. Similarly, if $\eta$ is a finite Borel partition of $X\times X$, then $\xi=\{\phi^{-1}(D\cap\Delta):D\in\eta\}$ is a finite partition of $\Delta=\{(x,x):x\in X\}$, such that ${\rm diam}(\xi)\leq{\rm diam}(\eta)$. Moreover, one has
$H_{\rho_{k}}(\xi^{n})=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=H_{\rho_{k}^{\ast}}(\eta^{n}),$
which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\geq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. This shows (4.5).
Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e. the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$. Hence
@@
This shows $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
@@
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}\geq\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Since $\gamma$ is arbitrary, by Proposition 2.1 one has
$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
The case of $\overline{\rm mdim}_{M}(f,X,d)=\infty$ is still valid by considering a slight modification.
For sufficiently small $\gamma>0$, let $c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}$. By Theorem 2.4, we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f,\xi)>c_{\gamma}.$ (4.6)
Put $\hat{K}=\overline{{\rm Conv}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}}$. Then the similar argument gives us
$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d),$
which implies the desired result by Proposition 2.1.
Finally, we give an example to verify the Corollary 4.2 and Theorem 4.3.
@@
Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric
@@
and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})\mapsto(x_{n+1})_{n\in\mathbb{Z}}$. It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
@@
    Thus, by the product formula of metric mean dimension one has ${\rm mdim}_{M}(\sigma\times\sigma,[0,1]^{\mathbb{Z}}\times[0,1]^{\mathbb{Z}},d\times d)=2$. Consequently, by Theorem 4.3, we have
    $\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
@@
    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, we have
@@
5 Examples
Applying the main results, we give the specific formulae of the metric mean dimensions of generic points of full shift over compact metric space.

### v3

@@ 4 Some applications of main results
In this section, we apply the main results to multifractal analysis of level sets and mean Li-Yorke chaos in infinite entropy systems. The main results in this section are presented in Theorems 4.1 and 4.3.
@@ 4.1 Variational principle of level sets
In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ within the context of an abstract framework.
Given a dynamical system $(X,f)$, assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function. We define
$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right].$
For any $\alpha\in L_{\varphi}$, we define $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$. The continuity of $\varphi$ then implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$. We define the level set of $\alpha$ w.r.t. $\varphi$ as
$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\},$
which consists of the set of points of $X$ such that $\varphi$ takes constant value $\alpha$ on $V_{f}(x)$.
Condition A: For every $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
Several examples of the function $\varphi$ that satisfy Condition A are stated below:
@@
    We define
    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu},$
    where $\psi$, $\phi\in C(X)$ and $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$).
@@
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, that is, for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
@@
    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for all $\alpha\in L_{\Phi}$. Specially, let $\psi\in C(X)$ and define $\phi_{n}=S_{n}\psi$, then $\Phi=\{\phi_{n}\}$ is an additive potential, in this case we have
@@
Given a nonempty open subset $U\subset X$, we similarly define the mixed $\alpha$-level set
@@
Let $(X,d,f)$ be a t.d.s. with the specification property, and let $U\subset X$ be a nonempty open subset. Assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A. Then for any $\alpha\in L_{\varphi}$ and $S\in\{B,P\}$,
@@
Define $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.17), there exists a constant $c>0$ such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$
holds for all $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$. Together with the fact that $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
@@
for all sufficiently small $\epsilon>0$. This implies that
@@
Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. Fix $x\in X$. Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, it follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
@@
A direct consequence of Theorem 4.1 is the variational principles for the Bowen metric mean dimension of level sets, as derived by Backes and Rodrigues [BR23, Theorem A]. However, we are able to strengthen this result by demonstrating that the variational principle remains valid for the packing metric mean dimension of level sets [^3] Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure..
@@
Let $(X,d,f)$ be a t.d.s. with the specification property and $\psi$ be a continuous function on $X$. Let
@@
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
@@
Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We obtain the desired result by Theorem 4.1. ∎
@@ 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In the present subsection, we establish a connection between the concepts of metric mean dimension and mean Li-Yorke chaos within the context of infinite entropy. We provide a quantitative description of the chaotic behavior exhibited by infinite entropy systems. Specifically, we demonstrate that the packing and upper capacity metric mean dimensions of the set of mean Li-Yorke pairs are equal to twice the metric mean dimension of the phase space, as stated in Theorem 4.3.
Given a t.d.s. $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric defined as
$d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\},$
and the mapping $f\times f:X\times X\rightarrow X\times X$, given by $(f\times f)(x,y):=(fx,fy)$, is continuous.
A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,~{}\text{and}\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
The set of all mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$. A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set. The well-known fact is that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]).
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Then for $S\in\{UC,P\}$ we have
$\overline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$
$\underline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~{}\underline{\rm mdim}_{M}(f,X,d).$
@@
It suffices to show the first equality since the second one is obtained similarly.
We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
Notice that the specification property implies that the system has positive topology entropy. Since positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that
$\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$
Set $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then, in the weak-topology any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
Now take $\nu\in M_{f}(X)$ arbitrarily and put $\mu_{1}:=\phi_{\ast}\nu=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. It is easy to check that $\mu_{1}\in M_{f\times f}(X\times X)$, and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. For sufficiently small $\gamma>0$, let
$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)-\gamma\}.$
By Theorem 2.4, we pick a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and a sequence of invariant measures $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$ (4.3)
Let $K=\overline{{\rm Conv}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}$, i.e., the convex closure of the set $\{\mu_{0},\mu_{1}\}\cup\{m_{k}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. It is easy to see that if two systems admit the specification property, so does the product system. By [PS07], we have $G_{K}\not=\emptyset$. For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}_{n\in\mathbb{N}}$ and $\{t_{n}^{1}\}_{n\in\mathbb{N}}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively. Hence,
@@
This implies that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
@@
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)}{|\log\epsilon_{k}|}\geq c_{\gamma}.$
Since $\gamma$ is arbitrary, one has
$\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
$\leq$ $\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)$
$\leq$ $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d),$
where we used the Proposition 2.1 for the second inequality. Together with $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$, we obtain the desired equality. ∎
Finally, we give an example to verify these results presented in Corollary 4.2 and Theorem 4.3.
@@
Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric:
@@
and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})=(x_{n+1})_{n\in\mathbb{Z}}$.
It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
@@
    Thus, by Theorem 4.3, we have
    $\overline{\rm mdim}_{M}^{P}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
@@
    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Corollary 4.2 we have
@@
4.3 Generic points of full shift over compact metric space
In this subsection, based on the main results, we present some specific formulae for the metric mean dimensions of the set of generic points of full shifts over compact metric spaces.
We begin by reviewing some definitions and results for our statements. Let $(X,d,f)$ be a t.d.s., $\mu\in M_{f}(X)$, $\epsilon>0$ and $x\in X$. We define the Brin-Katok local entropy at $x$ as
$\overline{h}_{\mu}^{BK}(x,\epsilon)=\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n},$
and the Brin-Katok local $\epsilon$-entropy of $\mu$ as
$\overline{h}_{\mu}^{BK}(f,\epsilon):=\int\overline{h}_{\mu}^{BK}(x,\epsilon)\mathop{}\!\mathrm{d}\mu,$
respectively. Specially, if $\mu\in M_{f}^{e}(X)$, then $h_{\mu}^{BK}(x,\epsilon)=h_{\mu}^{BK}(f,\epsilon)$ holds for $\mu$-a.e. $x\in X$.